The intensity I of light from a light bulb varies inversely as the square of the distance d from the light bulb. Suppose I is 1000 w/m^2 when the distance is 4m. How much farther would it be to a point where the intensity is 810 w/m^2?
Answer by Alan3354(69443)

Do it like this one:
---------
The intensity I of light from a light bulb varies inversely as the square of the distance d from the light bulb. Suppose I is 640 w/m^2 when the distance is 8m. How much farther would it be to a point where the intensity is 90 w/m^2?
-------------------
k is a constant, d = distance
640 = k/8^2 = k/64
k = 40960
------------
For 90 w/m^2:
90 = 40960/d^2
d^2 = 40960/90
d = 64/3 meters